it takes 1 person 4500 minutes to take down the seats, how many people will they need to do it in 6 hours?
Answer by Theo(13342)

rate * people * time = quantity of work.

the rate is the average rate per person.

you get r*p*t = q

when you have 1 person doing the work, it takes 4500 minutes.

you get p = 1 and t = 4500 in the equation.

the formula becomes:

r*1*4500 = q

solve for r to get r = q/4500.

you want to know how many people it takes to do the same job in 6 hours.

the rate that each person works is presumably the same, which is q/4500.

6 hours is equal to 6 * 60 = 360 minutes.

r is equal to q/4500
t is equal to 360
p is what you want to solve for.

the equation of r*p*t = q becomes:

q/4500 * p * 360 = q

multiply both sides of the equation by 4500 / q to get p * 360 = q * 4500 / q.

simplify to get p * 360 = 4500

divide both sides of this equation by 360 to get p = 4500 / 360.

this results in p = 12.5.

if 1 person does it in 4500 minutes, then 12.5 people are required to do it in 360 minutes.

the value of q was irrelevant because it canceled out.

this means you could have replaced q with any number and the results would be the same.